OpenStudy (lilshane):

Betty swam 100 meters in 2 minutes. Which rate represents this situation? A. B. C. D.

OpenStudy (lilshane):

a \[a \frac{ 100 mins }{ 2 mins }\]

OpenStudy (lilshane):

\[b \frac{ 100 meters }{ 2 meters }\]

OpenStudy (lilshane):

\[c \frac{ 100 mins }{ 2 meters }\]

OpenStudy (lilshane):

d \[d \frac{ 100 meters }{ 2 mins }\]

OpenStudy (lilshane):

@AnswerMyQuestions

OpenStudy (lilshane):

@iGreen

OpenStudy (lilshane):

@jordan123321

OpenStudy (igreen):

Write 100 meters being divided by 2 minutes.

OpenStudy (lilshane):

so thts 50 meters in 1 minutes

OpenStudy (lilshane):

i think my answer is d

OpenStudy (igreen):

Yep, you got it.

OpenStudy (lilshane):

can you help me with some more